OpenStudy (anonymous):

When 15 g of NaCl is dissolved in water to a final volume of 500 mL the concentration of the solution is 500 mM. What is the molecular weight of NaCl?

OpenStudy (aaronq):

sub in: \(\sf moles =\dfrac{mass}{Molar~mass}\) into: \(\sf Molarity=\dfrac{moles}{L_{solution}}\) \(\rightarrow \sf Molarity=\dfrac{\dfrac{mass}{Molar~mass}}{L_{solution}}=\dfrac{mass}{Molar~ mass*L_{solution}}\) solve for molar mass